# An Extended Closure Relation by LightGBM for Neutrino Radiation Transport in Core-collapse Supernovae

### v1

@@ 3.2 Feature Engineering
The closure relation is originally meant to give high-order moments such as the Eddington tensor in terms of the lower-order moments like the energy density and flux. In our previous attempt to build a deep neural network to do this task, we realized that the lower-order moments are just not sufficient and added some other local features such as the matter velocity and its shear to the input. In this paper, we extend this idea. Thanks to the efficiency of LightGBM, we are able to incorporate much more information, some of them nonlocal. It is true that the more model-specific information we employ, the less general the machine becomes, but the purpose of this paper is to construct a machine learning model that we can apply to the CCSN simulation with the truncated moment method.
@@
In providing a large number of features to the machine, we find feature engineering very important. By feature engineering, we mean finding better combinations of the original features, normalizing features and, if necessary, removing the basic features after using them for feature engineering. In principle, this process does not increase the information included in the original features. If the output depends on a function of rather involved combinations of these features, the machine may need a long time or even fail to train itself even if those features contain sufficient information. On the other hand, if the human knows a priori those combinations of the original features and teaches them to the machine, we can expect the training efficiency to be improved substantially. Unfortunately, we do not know what should be the best combinations. Thanks to the efficiency of LightGBM, we can try various combinations. The engineered features given below are indeed obtained that way.
@@
As mentioned above, we will incorporate nonlocal quantities to the input features. This is probably understandable if one recalls the fact that the radiative transfer is most difficult to treat in the semi-transparent region, where the mean free path of neutrino becomes comparable to the scale height of matter and the neutrino distribution is not determined locally indeed. It is noted that the basic features in Table 1 include the optical depth, which is nonlocal information. As explained below, we add more.
The engineered features thus produced are classified into three types: advanced features, spatial-shift features, and spatial-difference features. The advanced features, listed in Table 2, are those features given as functions of the basic features. Particularly important for the improvement of the model are the flux factor in the LB frame $\chi_{\rm LB}=|F_{\rm LB}^{i}|/E_{\rm LB}$, the colinearity parameter $\frac{F_{\rm{LB}}^{j}v_{j}}{E_{\rm{LB}}|v^{i}|}$ that measures the alignment of the neutrino flux and matter velocity, and the parameter we call the“neutrinosphere indicator,” which is a binary indicator of whether the point is well inside the neutrinosphere or not, i.e., it is 1 if the optical depth $\tau$ is greater than 5 and 0 otherwise.
Table 2: Advanced Features
Feature Definition Dimension
Neutrino energy $\varepsilon_{\nu}$ (MeV)
Flux factor in LB frame $\chi_{\rm LB}=|F_{\rm LB}^{i}|/E_{\rm LB}$
Colinearity parameter $\frac{F_{\rm{LB}}^{j}v_{j}}{E_{\rm{LB}}|v^{i}|}$
Sign of $v^{1}$ in FP frame $I(v_{\text{FP}}^{1}\geq 0)$a -
Sign of $v^{2}$ in FP frame $I(v_{\text{FP}}^{2}\geq 0)$
In neutron sphere $I(\tau\geq 5)$
Logarithm of flux factor $\log(\chi_{\text{LB}})$
Logarithm of optical depth $\log(\tau)$
Logarithm of baryon mass density $\log(\rho_{\text{baryon}})$ $\log(\text{g/cm}^{3})$
@@
The spatial-shift features are the features concerning the neighboring grids. The closure relation is normally local, i.e., it is a relation among physical quantities at the same location. It turns out, however, that the inclusion of the neighboring physical quantities as input improves the accuracy of the inference. More concretely, we incorporate the flux factor $\chi_{\rm LB}=|F_{\rm LB}^{i}|/E_{\rm LB}$ and the colinearity parameter $\frac{F_{\rm{LB}}^{j}v_{j}}{E_{\rm{LB}}|v^{i}|}$ not only on the grid point of concern but also on the neighboring grid points. For instance, let $Q_{k,l}$ represent one of the aforementioned quantities at the $k$-th radial grid point and the $l$-th angular grid point; we then incorporate $Q_{k-5i,l}$ (where $i=-1,1,2,\dots,6$), $Q_{k+1,l}$, $Q_{k,l+1}$, $Q_{k-1,l}$ and $Q_{k,l-1}$ as components of the input feature vector for $Q_{k,l}$. [^4] If it goes beyond the spatial boundaries, those points are simply ignored. In particular, it was found that incorporating information from the inner regions of the star, such as $Q_{k-5i,l}$ (where $i=1,2,\dots,6$) and $Q_{k-1,l}$, significantly improves the prediction accuracy. The reasoning here is that since neutrinos predominantly move outward, upstream information from radially inner points is particularly valuable.
We also find it useful to employ the difference between the feature on the grid of concern and that on the inner adjacent grid:
$\Delta Q_{k,\ell}=Q_{k,\ell}-Q_{k-1,\ell}.$ (11)
to which we refer as the spatial-difference feature. Although it may seem redundant, we include both the spatial-shift and difference features in the input as we do not know which feature is more effective. Note that irrelevant features are automatically ignored in LightGBM.
@@ 3.3 Models
The pipeline of the model employed in this study is shown in Figure 2. After creating new features from the basic features through feature engineering, we conduct training with the 8-fold cross validation (CV), in which the data set is randomly divided into 8 subsets to build 8 models. One of them serves as the validation set and the remaining 7 as the training data. They make predictions individually and the arithmetic mean of those predictions is adopted as the final prediction of this model. Such an approach is known to effectively prevent the model from over-fitting a particular dataset, thereby improving generalization.
@@
The main hyperparameters and their values in the LightGBM model are listed in Table 3. We employ the early stopping training, i.e., the training is terminated once the mean squared error is no longer lowered remarkably. It is also useful to avoid overfitting. As a consequence, we can set a sufficiently large value to n_estimators, the number of decision trees to be built. Though it is not our main goal in this paper to fine-tune the hyperparameters, we set different values to some hyperparameters for the diagonal components of the Eddington tensor from those for the off-diagonal components as given in the table.
@@ 4 Results
In this paper, the data extracted from the snapshot at 300 ms after the bounce are used as the training set. The target variable is the Eddington tensor in the FP frame $k_{\rm{FP}}^{ij}$. We focus on the estimation of its 11-, 22-, 33- and 12-components, since they are the diagonal and main off-diagonal elements, respectively. The test data were collected from the snapshots at 100 ms, 150 ms, 200 ms and 250 ms after the bounce. In the following, the diagonal and off-diagonal components are considered separately in turn.
@@ 4.1 Diagonal Components
The diagonal components of the Eddington tensor play the main role in neutrino transport, and hence their accurate estimation is crucially important. The conventional M1-closure relation does not reproduce it perfectly (Nagakura et al., 2018; Harada et al., 2019, 2020; Iwakami et al., 2022), and the machine-learning closure is expected to provide better predictions.
Figure 3 shows the learning curve in the training of the 11-component of the Eddington tensor. The training is conducted to minimize the Mean Squared Error (MSE):
$L_{\varepsilon}^{(2)}=\frac{1}{|D_{\varepsilon}|}\sum_{D_{\varepsilon}}|k_{\rm boltz,FP}^{ij}-k_{\rm infer,FP}^{ij}|^{2},$ (12)
where the sum is taken over the subset $D_{\varepsilon}$ of the data corresponding to a particular neutrino energy and $|D_{\varepsilon}|$ is the number of its elements; $k_{\rm boltz,FP}^{ij}$ is the Eddington tensor obtained by the Boltzmann-radiation-hydrodynamics simulation whereas $k_{\rm infer,FP}^{ij}$ is the one obtained with our LightGBM model. As the solid lines show, the training is successful with the values of MSE decreasing monotonically down to $<10^{-5}$ for all 8 models. These lines are terminated at different numbers of estimator because of the early stopping. The evaluation of the validation data also presented in the figure as dashed lines indicates that the overfitting is avoided reasonably well indeed. Similar results are obtained for other components of the Eddington tensor.
Figure 3: Each plot represents the loss for each fold in cross-validation. The vertical axis shows the mean squared error, while the horizontal axis represents the number of estimators.
We move on to the predictions of these models. In Figures 4, 5, 6 we plot the mean absolute error (MAE) for the diagonal components of the Eddington tensor for different post-bounce times: 250, 200, 150, and 100 ms. MAE is defined as
@@
where $k_{\rm infer,FP}^{ij}$ is the $ij$-component of the Eddington tensor obtained either with our Light-GBM model or via the M1-closure relation. It is apparent that our LightGBM closure gives more accurate predictions than the M1 closure for almost all energies at these times. The behavior of MAE is essentially the same among the three diagonal components. As the time goes back to earlier times, the accuracy is degraded somewhat as expected. It is encouraging, however, that the results are still better than those for the M1 closure except at $\sim$ 30 MeV, where the M1 closure gives exceptionally good results.
@@
Figure 4: Relationship between neutrino energy and prediction error. The target variable is $k_{\rm{FP}}^{11}$ and the validation data correspond to 250ms (top left), 200ms (top right), 150ms (bottom left) and 100ms (bottom right) after the core bounce. The green and red lines represent the MAE obtained using the M1-closure relation and our LightGBM prediction, respectively.
Figure 5: The same as figure 4 except that the target variable is $k_{\rm FP}^{22}$.
Figure 6: The same as figure 4 except that the target variable is $k_{\rm FP}^{33}$.
@@
Although MAE is useful to see the overall convergence of the model to the target, what is more important is to what extent the individual components of the Eddington tensor are reproduced at different points in space. We hence compare next the radial profiles of the components of the Eddington tensors along the radial line at $\theta=\pi/2$ in figures 718. The results are not much different for other angles.
Figure 7: Comparison of the radial profiles of $k_{\rm FP}^{11}$ along $\theta=\pi/2$. The horizontal axis represents the radius, while the vertical axis indicates the Eddington tensor. The blue, green, and red lines correspond to the Boltzmann-radiation-hydrodynamics simulations, the M1 closure relation, and the prediction obtained by LightGBM, respectively.Different panels correspond to different neutrino energies: $1.0\,{\rm MeV}$ (upper left), $8.5\,{\rm MeV}$ (upper right), $18.8\,{\rm MeV}$ (lower left), and $54.0\,{\rm MeV}$ (lower right).
Figure 8: The same as figure 7 except that the snapshot time is $200\,{\rm ms}$ and the target is $k_{\rm{FP}}^{11}$.
Figure 9: The same as figure 7 except that the snapshot time is $150\,{\rm ms}$ and the target is $k_{\rm{FP}}^{11}$.
Figure 10: The same as figure 7 except that the snapshot time is $100\,{\rm ms}$ and the target is $k_{\rm{FP}}^{11}$.
Figure 11: The same as figure 7 except that the snapshot time is $250\,{\rm ms}$ and the target is $k_{\rm{FP}}^{22}$.
Figure 12: The same as figure 7 except that the snapshot time is $200\,{\rm ms}$ and the target is $k_{\rm{FP}}^{22}$.
Figure 13: The same as figure 7 except that the snapshot time is $150\,{\rm ms}$ and the target is $k_{\rm{FP}}^{22}$.
Figure 14: The same as figure 7 except that the snapshot time is $100\,{\rm ms}$ and the target is $k_{\rm{FP}}^{22}$.
Figure 15: The same as figure 7 except that the snapshot time is $250\,{\rm ms}$ and the target is $k_{\rm{FP}}^{33}$.
Figure 16: The same as figure 7 except that the snapshot time is $200\,{\rm ms}$ and the target is $k_{\rm{FP}}^{33}$.
Figure 17: The same as figure 7 except that the snapshot time is $150\,{\rm ms}$ and the target is $k_{\rm{FP}}^{33}$.
Figure 18: The same as figure 7 except that the snapshot time is $100\,{\rm ms}$ and the target is $k_{\rm{FP}}^{33}$.
@@
We can confirm again that the LightGBM closure outperforms the M1 closure in general. Of particular note is that our LightGBM model can reproduce the values of the 11-component of the Eddington tensor smaller than 1/3 and the 22- and 33-components greater than 1/3 at rather small radii. Those values are obtained when the opacity decreases rapidly with the radius in the vicinity of the neutrinosphere; then the neutrino distribution is almost isotropic in the outward hemisphere whereas it is heavily depleted in the opposite hemisphere (see the discussion in Harada et al., 2022). By its construction the M1 closure cannot treat such angular distributions properly. It is noted that our LightGBM fairs even better than the TBNN model in Harada et al. (2022).
@@
It is also worth mentioning that the LightGBM model can handle the shock wave with reasonable accuracies. The shock wave is located at $170\,{\rm km}$, $140\,{\rm km}$, $130\,{\rm km}$, $120\,{\rm km}$, and $110\,{\rm km}$ at $100\,{\rm ms}$, $150\,{\rm ms}$, $200\,{\rm ms}$, $250\,{\rm ms}$, and $300\,{\rm ms}$ post bounce, respectively. The Eddington tensor is discontinuous at the shock wave. The jump gets larger as the energy increases. The radial profiles of the Eddington tensor for the energy of $54\,{\rm MeV}$ presented in figures 718 demonstrate indeed that the predictions of the LightGBM model are closer to the results of the Boltzmann simulations than the M1-closure relation in general.
@@
However, the Eddington tensor predicted by LightGBM has several issues. Firstly, the predicted values are not so smooth as those obtained with the M1 closure. This tendency becomes more pronounced as we go back in time and see the predictions themselves degraded. In fact, the Eddington tensor at 250 ms post bounce plotted in Figures 7, 11, 15 is relatively smooth, whereas the lines are more jagged at 100 ms after bounce as shown in Figures 10, 14, 18. We think that this phenomenon occurs because the current predictions are made for each grid point individually and that the issue could be resolved by adopting a machine learning model capable of incorporating non-local information into the predictions. Machine learning models such as Convolutional Neural Networks (CNNs) (LeCun et al., 1998) or Recurrent Neural Networks (RNNs) (Hochreiter & Schmidhuber, 1997) may have robustness to local noises and outliers in the input data, thereby giving smoother predictions.
@@
It should be also mentioned that the current model does not take into account physical requirements. For instance, the trace of the Eddington tensor is unity by definition. There is no guarantee, however, that the predicted results satisfy this constraint although the violation is quite minor in the current models, as can be seen in Figure 19. This issue may be addressed with physics-informed neural networks (Raissi et al., 2019): for example, one may incorporate the violations of constraints into the loss function so that they could be minimized. Although these possibilities are interesting on their own rights and warrant further investigations, they are much beyond the scope of this paper and will be future works.

### v2

@@ 3.2 Feature Engineering
The closure relation is originally meant to give high-order moments such as the Eddington tensor in terms of lower-order moments like the energy density and flux. In our previous attempt to build a deep neural network to do this task, we realized that the lower-order moments are just not sufficient and added some other local features such as the matter velocity and its shear to the input. In this paper, we extend this idea. Thanks to the efficiency of LightGBM, we are able to incorporate much more information, some of which are non-local. It is true that the more model-specific information we employ, the less general the machine becomes, but the purpose of this paper is to construct a machine learning model that we can apply to the CCSN simulation with the truncated moment method.
@@
In providing a large number of features to the machine, we find feature engineering very important. By feature engineering, we mean finding better combinations of the original features, normalizing features, and, if necessary, removing the basic features after using them for feature engineering. In principle, this process does not increase the information included in the original features. If the output depends on a function of rather involved combinations of these features, the machine may need a long time or even fail to train itself even if those features contain sufficient information. On the other hand, if the human knows a priori those combinations of the original features and teaches them to the machine, we can expect the training efficiency to be improved substantially. Unfortunately, we do not know what the best combinations should be. Thanks to the efficiency of LightGBM, we can try various combinations. The engineered features given below are indeed obtained that way.
@@
As mentioned above, we will incorporate non-local quantities to the input features. This is probably understandable if one recalls the fact that the radiative transfer is most difficult to treat in the semi-transparent region, where the mean free path of neutrinos becomes comparable to the scale height of matter, and the neutrino distribution is indeed not determined locally.
The engineered features thus produced are classified into three types: combined features, non-local features, and differential features.
Table 1: The features employed in the model
Class Feature Definition Dimension Adopteda
basic Neutrino energy $\varepsilon$ $\rm{(MeV)}$ D/OD
Optical depth $\tau$ OD
Baryon mass density $\rho_{\text{baryon}}$ $\rm{(g/cm^{3})}$ OD
Energy density (LB) $E_{\text{LB}}$ $\rm{(erg/cm^{3})}$ OD
Flux (LB) $F_{\text{LB}}^{i}\,(i=r,\theta,\phi)$ $\rm{(erg/cm^{3})}$ OD
Fluid velocity (LB) $v^{i}\,(i=r,\theta,\phi)$ $\rm{(cm/s)}$ OD
combined Flux factor (LB) $\chi_{\rm LB}=|F_{\rm LB}|/E_{\rm LB}$ D/OD
collinearity parameter $F_{\rm{LB}}^{k}v_{k}/E_{\rm{LB}}|v|$ D/OD
Sign of $v^{j}\;(j=1,2)$ (FP) $I(v_{\text{FP}}^{j}\geq 0)$b $\;(j=1,2)$ D/OD
In neutrinosphere $I(\tau\geq 5)$ D/OD
Logarithm of flux factor $\log(\chi_{\text{LB}})$ D/OD
Logarithm of optical depth $\log(\tau)$ D/OD
Logarithm of baryon mass density $\log(\rho_{\text{baryon}}/({\rm g/cm^{3}}))$ D/OD
Fluid speed $|v|$ (cm/s) OD
Fluid velocity (FP) $v_{\rm FP}^{i}\;(i=1,2,3)$ (cm/s) OD
Normalized fluid velocity (FP) $v_{\rm FP}^{i}/|v|\;(i=1,2,3)$ OD
Neutrino mean velocity (FP) $F_{\rm FP}^{i}/E\;(i=1,2,3)$ OD
near non-local Flux factor (LB) $\chi_{\rm LB}=|F_{\rm LB}|/E_{\rm LB}$ D/OD
collinearity parameter $F_{\rm{LB}}^{k}v_{k}/E_{\rm{LB}}|v|$ D/OD
Fluid velocity (FP) $v_{\rm FP}^{i}\;(i=1,2)$ (cm/s) OD
Normalized fluid velocity (FP) $v_{\rm FP}^{i}/|v|\;(i=1,2)$ OD
Fluid speed $|v|$ (cm/s) OD
Neutrino mean velocity (FP) $F_{\rm FP}^{i}/E\;(i=1,2)$ OD
far non-local Flux factor (LB) $\chi_{\rm LB}=|F_{\rm LB}|/E_{\rm LB}$ D/OD
collinearity parameter $F_{\rm{LB}}^{k}v_{k}/E_{\rm{LB}}|v|$ D/OD
differential Flux factor (LB) $\partial_{j}\chi_{\rm LB}\;(j=r,\theta)$ ($\rm cm^{-1}$) or - D/OD
collinearity parameter $\partial_{j}(F_{\rm{LB}}^{k}v_{k}/E_{\rm{LB}}|v|)\;(j=r,\theta)$ ($\rm cm^{-1}$) or - D/OD
Fluid velocity (FP) $\partial_{j}v_{\rm FP}^{i}\;(i=1,2;\;j=r,\theta)$ ($\rm s^{-1}$) or ($\rm cm/s$) OD
Normalized fluid velocity (FP) $\partial_{j}(v_{\rm FP}^{i}/|v|)\;(i=1,2;\;j=r,\theta)$ ($\rm cm^{-1}$) or - OD
Fluid speed $\partial_{j}|v|\;(j=r,\theta)$ ($\rm s^{-1}$) or ($\rm cm/s$) OD
Neutrino mean velocity (FP) $\partial_{j}(F_{\rm FP}^{i}/E)\;(i=1,2;\;j=r,\theta)$ ($\rm cm^{-1}$) or - OD
@@
    “Adopted” column indicates whether each feature is adopted in both diagonal and off-diagonal prediction (O/OD) or in only off-diagonal prediction (OD).
  • b
@@
The combined features considered here is listed as ll. 7–17 in Table 1. Particularly important for the improvement of the model are the flux factor in the LB frame $\chi_{\rm LB}=|F_{\rm LB}^{i}|/E_{\rm LB}$, the collinearity parameter $\frac{F_{\rm{LB}}^{k}v_{k}}{E_{\rm{LB}}|v^{i}|}$ that measures the alignment of the neutrino flux and matter velocity, and the parameter we call the ”neutrinosphere indicator” which is a binary indicator of whether the point is well inside the neutrinosphere or not, i.e., it is 1 if the optical depth $\tau$ is greater than 5 and 0 otherwise.
The non-local features are the features concerning the neighboring grid points. The closure relation is normally local; i.e., it is a relation among physical quantities at the same location. It turns out, however, that the inclusion of the neighboring physical quantities as input improves the accuracy of the inference. More concretely, we incorporate the features listed as ll. 18–23 in Table 1 not only on the grid point of concern but also on the neighboring grid points. Here, we further classify the non-local features into near non-local and far non-local ones. Let $Q_{k,l}$ represent one of the listed features at the $k$-th radial grid point and the $l$-th angular grid point; we then incorporate $Q_{k-1,l}$, $Q_{k+1,l}$, $Q_{k,l-1}$, and $Q_{k,l+1}$ as near non-local features (ll. 18–21 in the table), and $Q_{k-1,l}$, $Q_{k+1,l}$, $Q_{k,l-1}$, $Q_{k,l+1}$, and $Q_{k-5i,l}$ ($i=-1,1,2,\dots,6$) as far non-local features (ll. 22–23 in the table) [^4] If it goes beyond the spatial boundaries, those points are simply ignored.. We find in particular that the incorporation of the inner points, $Q_{k-5i,l}$ (where $i=1,2,\dots,6$) and $Q_{k-1,l}$, significantly improves the prediction accuracy. The reasoning here is that since neutrinos predominantly move outward, the upstream information, i.e., the information on the radially inner points is particularly valuable.
We also find it useful to employ the derivatives of features :
$\frac{\Delta Q_{k,\ell}}{\Delta r}=\frac{Q_{k,\ell}-Q_{k-1,\ell}}{r_{k,\ell}-r_{k-1,\ell}}.$ (11)
$\frac{\Delta Q_{k,\ell}}{\Delta\theta}=\frac{Q_{k,\ell}-Q_{k,\ell-1}}{r_{k,\ell}(\theta_{k,\ell}-\theta_{k,\ell-1})}.$ (12)
to which we refer as the differential features. Although it may seem redundant, we include both the non-local and differential features in the input as we do not know which one is more effective. Note that irrelevant features are automatically ignored in LightGBM.
@@ 3.3 Models
The pipeline of the model employed in this study is shown in Figure 2. After creating new features from the basic features through feature engineering, we conduct training with the 8-fold cross-validation (CV), in which the data set is randomly divided into 8 subsets to build 8 models. One of them serves as the validation set and the remaining 7 as the training data. They make predictions individually and the arithmetic mean of those predictions is adopted as the final prediction of this model. Such an approach is known to effectively prevent the model from over-fitting a particular dataset, thereby improving generalization.
@@
The main hyperparameters and their values in the LightGBM model are listed in Table 2. We employ the early stopping training, i.e., the training is terminated once the mean squared error is no longer lowered remarkably. It is also useful to avoid overfitting. As a consequence, we can set a sufficiently large value to n_estimators, the number of decision trees to be built. Though it is not our main goal in this paper to fine-tune the hyperparameters, we set some hyperparameters separately for the diagonal components of the Eddington tensor and for the off-diagonal components.
@@ 4 Results
In this paper, the data extracted from the snapshot at 300 ms after bounce is used as the training set. The target variable is the Eddington tensor in the FP frame $k_{\rm{FP}}^{ij}$. We focus on the estimation of its 11-, 22-, 33-, and 12-components, since they are the diagonal and main off-diagonal elements, respectively. The test data were collected from the snapshots at 100 ms, 150 ms, 200 ms, and 250 ms after bounce. In the following, the diagonal and off-diagonal components are considered separately in turn.
@@ 4.1 Diagonal Components
The diagonal components of the Eddington tensor play the main role in neutrino transport, and hence their accurate estimation is crucially important. The conventional M1 closure relation does not reproduce it perfectly (Nagakura et al., 2018; Harada et al., 2019, 2020; Iwakami et al., 2022), and the machine-learning closure is expected to provide better predictions.
The training is conducted to minimize the Mean Squared Error (MSE):
$L^{(2)}=\frac{1}{|D|}\sum_{D}|k_{\rm boltz,FP}^{ij}-k_{\rm infer,FP}^{ij}|^{2},$ (13)
where the sum is taken over the data and $|D|$ is the number of its elements; $k_{\rm boltz,FP}^{ij}$ is the Eddington tensor obtained by the Boltzmann-radiation-hydrodynamics simulation whereas $k_{\rm infer,FP}^{ij}$ is the one obtained with our LightGBM model.
We move on to the predictions of these models. In Figure 3 we plot the mean absolute error (MAE) for the diagonal components of the Eddington tensor for different post-bounce times: 250, 200, 150, and 100 ms. MAE is defined as
@@
where $k_{\rm infer,FP}^{ij}$ is the $ij$-component of the Eddington tensor obtained either with our LightGBM model or via the M1 closure relation. It is apparent that our LightGBM closure gives more accurate predictions than the M1 closure for almost all energies at these times. The behavior of MAE is essentially the same among the three diagonal components. As the time goes back to earlier times, the accuracy is degraded somewhat as expected. It is encouraging, however, that the results are still better than those for the M1 closure except at $\sim$ 30 MeV, where the M1 closure gives exceptionally good results.
@@
Figure 3: Relationship between neutrino energy and prediction error. Each row corresponds to the plots of $k_{\rm{FP}}^{11}$, $k_{\rm{FP}}^{22}$, $k_{\rm{FP}}^{33}$, and $k_{\rm{FP}}^{12}$ from top to bottom, while each column represents the plots for 250 ms, 200 ms, 150 ms, and 100 ms after the core bounce from left to right. The green and red lines represent the MAE obtained using the M1 closure relation and our LightGBM prediction, respectively.
@@
Although MAE is useful to see the overall convergence of the model to the target, what is more important is to what extent the individual components of the Eddington tensor are reproduced at different points in space. We hence compare next the radial profiles of the components of the Eddington tensors along the radial line at $\theta=\pi/2$ in figures 46. The results are not much different for other angles. At the bottom of each plot, we show the absolute error, i.e., the magnitude of deviation from the result of the Boltzmann simulation.
[…]
Figure 5: The same as figure 4 except that the target is $k_{\rm{FP}}^{22}$.
Figure 6: The same as figure 4 except that the target is $k_{\rm{FP}}^{33}$.
Figure 7: The same as figure 4 except that the target is $k_{\rm{FP}}^{12}$.
@@
We can confirm again that the LightGBM closure outperforms the M1 closure in general. Of particular note is that our LightGBM model can reproduce the values of the 11-component of the Eddington tensor smaller than 1/3 and the 22- and 33-components greater than 1/3 at rather small radii. Those values are obtained when the opacity decreases rapidly with the radius in the vicinity of the neutrinosphere; then the neutrino distribution is almost isotropic in the outward hemisphere whereas it is heavily depleted in the opposite hemisphere (see the discussion in Harada et al., 2022). By its construction, the M1 closure cannot treat such angular distributions properly. It is noted that our LightGBM fares even better than the TBNN model in Harada et al. (2022).
@@
It is also worth mentioning that the LightGBM model can handle the shock wave with reasonable accuracy. The shock wave is located at $170\,{\rm km}$, $140\,{\rm km}$, $130\,{\rm km}$, $120\,{\rm km}$, and $110\,{\rm km}$ at $100\,{\rm ms}$, $150\,{\rm ms}$, $200\,{\rm ms}$, $250\,{\rm ms}$, and $300\,{\rm ms}$ post-bounce, respectively. The Eddington tensor is discontinuous at the shock wave. The jump gets larger as the energy increases. The radial profiles of the Eddington tensor for the energy of $54\,{\rm MeV}$ presented in Figures 46 demonstrate indeed that the predictions of the LightGBM model are closer to the results of the Boltzmann simulations than the M1 closure relation in general.
It should be noted, however, that the feature importance represents the degree to which the elimination of a single feature degrades the accuracy of prediction. The metric presented here may not fully capture the importance of each feature. It may not reflect its true importance even qualitatively when multiple features work collectively in an intricate way. We hence need caution to interpret the result.
Features Feature Importance
$\chi(r_{i},\theta_{j})$ $4.5\times 10^{5}$
$\log\chi(r_{i},\theta_{j})$ $3.0\times 10^{5}$
$\chi(r_{i+1},\theta_{j})$ $3.5\times 10^{4}$
$\log\rho_{\mathrm{baryon}}(r_{i},\theta_{j})$ $8.7\times 10^{3}$
$\chi(r_{i-5},\theta_{j})$ $7.0\times 10^{3}$
$\chi(r_{i-1},\theta_{j})$ $6.9\times 10^{3}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-10},\theta_{j})$ $2.5\times 10^{3}$
$\epsilon_{\nu}$ $2.0\times 10^{3}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-5},\theta_{j})$ $9.8\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i},\theta_{j})$ $9.3\times 10^{2}$
$\log\tau(r_{i},\theta_{j})$ $9.1\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-1},\theta_{j})$ $8.4\times 10^{2}$
$\frac{\Delta\chi}{\Delta r}(r_{i},\theta_{j})$ $7.0\times 10^{2}$
$\chi(r_{i-10},\theta_{j})$ $6.1\times 10^{2}$
$\chi(r_{i-5},\theta_{j})$ $5.3\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i+1},\theta_{j})$ $4.2\times 10^{2}$
$\chi(r_{i-30},\theta_{j})$ $3.9\times 10^{2}$
$\chi(r_{i-1},\theta_{j})$ $3.4\times 10^{2}$
$\chi(r_{i-15},\theta_{j})$ $2.1\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-15},\theta_{j})$ $1.6\times 10^{2}$
$\chi(r_{i},\theta_{j+1})$ $1.6\times 10^{2}$
$\chi(r_{i-25},\theta_{j})$ $1.5\times 10^{2}$
$\Delta\Bigl{(}\frac{F^{k}v_{k}}{E|v|}\Bigr{)}/\Delta\theta(r_{i},\theta_{j})$ $1.5\times 10^{2}$
$\chi(r_{i-20},\theta_{j})$ $1.3\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-25},\theta_{j})$ $1.2\times 10^{2}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i+1},\theta_{j})$ $1.2\times 10^{2}$
$\Delta\Bigl{(}\frac{F^{k}v_{k}}{E|v|}\Bigr{)}/\Delta r(r_{i},\theta_{j})$ $9.9\times 10^{1}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i+5},\theta_{j})$ $9.1\times 10^{1}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-1},\theta_{j})$ $7.4\times 10^{1}$
$\frac{\Delta\chi}{\Delta\theta}(r_{i},\theta_{j})$ $7.1\times 10^{1}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-20},\theta_{j})$ $6.9\times 10^{1}$
$\frac{F^{k}v_{k}}{E|v|}(r_{i-30},\theta_{j})$ $6.8\times 10^{1}$
$I(\tau\geq 5)(r_{i},\theta_{j})$ $2.3\times 10^{1}$
$\text{sign}(v^{2})(r_{i},\theta_{j})$ $2.0\times 10^{1}$
$\text{sign}(v^{1})(r_{i},\theta_{j})$ $3.2\times 10^{-1}$
Table 3: Feature importance for each feature in the model predicting the 11-component of the Eddington tensor (values rounded to 2 significant figures). The subscript LB for the flux factor and the collinearity parameter are omitted for simplicity.
@@
The Eddington tensor predicted by LightGBM has several issues. Firstly, the predicted values are not as smooth as those obtained with the M1 closure. This tendency becomes more pronounced as we go back in time and see the predictions themselves degraded. In fact, the Eddington tensor at 250 ms post-bounce plotted in Figures 46 is relatively smooth, whereas the lines are more jagged at 100 ms after bounce as shown in the same figures. We think that this phenomenon occurs because the current predictions are made for each grid point individually and that the issue could be resolved by adopting a machine learning model capable of incorporating non-local information into the predictions. Machine learning models such as Convolutional Neural Networks (CNNs) (LeCun et al., 1998) or Long Short Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) may have robustness to local noises and outliers in the input data, thereby giving smoother predictions.
@@
It should be also mentioned that the current model does not take into account physical requirements. For instance, the trace of the Eddington tensor is unity by definition. There is no guarantee, however, that the predicted results satisfy this constraint although the violation is quite minor in the current models, as can be seen in Figure 8. This issue may be addressed with physics-informed neural networks (Raissi et al., 2019): for example, one may incorporate the violations of constraints into the loss function so that they could be minimized. Although these possibilities are interesting in their own right and warrant further investigation, they are much beyond the scope of this paper and will be addressed in future work.